# Characterisation of distal actions of automorphisms on the space of one-parameter subgroups of Lie groups

### v2

@@ 4. Proofs of the main results
Step 6: Let $\psi,S:H\to H$ be such that $S|_{K}={\rm Id}$, $S|_{C(T)}=\alpha\,{\rm Id}$, $\psi|_{K}={\rm Id}$ and $\psi|_{C(T)}=\phi$, where $\alpha$ and $\phi$ are as in Step 2. Now $T|_{H}=S\psi=\psi S$, and $\psi$ is contained in a compact subgroup of ${\rm Aut}(H)$. Therefore, $S$ acts distally on ${\rm Sub}^{p}_{H}$ (cf. [41], Lemma 2.2), and $C(S)=C(T)$.
@@
Let $\{z_{t}=x_{t}c_{t}\}$ be a closed one-parameter subgroup, where $\{x_{t}\}\subset K$ is isomorphic to ${\mathbb{S}}^{1}$ and $\{c_{t}\}\subset C(S)=C(T)$ is a non-trivial (closed) one-parameter subgroup such that $S(c_{t})=c_{\alpha t}$, $t\in{\mathbb{R}}$ (as addition is the group operation on ${\mathbb{R}}^{d}$, we choose $c_{t}=tc$, $t\in{\mathbb{R}}$, for some nonzero $c\in{\mathbb{R}}^{d}$). Let $\{n_{k}\}$ be an unbounded sequence in ${\mathbb{N}}$ such that $S^{n_{k}}(\{z_{t}\})\to C$ for some $C$ in ${\rm Sub}^{p}_{H}$. Note that $\{z_{t}\}\subset\{x_{t}\}\times\{c_{t}\}$ and the latter is $S$-invariant and isomorphic to ${\mathbb{S}}^{1}\times{\mathbb{R}}$. So $C\subset\{x_{t}\}\times\{c_{t}\}$. For a fixed $t\in{\mathbb{R}}$, as $c_{t}\in C(S)=C(T)$, we get that
@@
As $\{x_{t}\}\times\{c_{t}\}$ is a direct product and $\{c_{t}\}$ is isomorphic to ${\mathbb{R}}$, it follows that $\beta=\alpha$ and $x_{t}=x_{\alpha t}$, i.e. $x_{(1-\alpha)t}=e$ for all $t\in{\mathbb{R}}$. As $\alpha\neq 1$, $x_{t}=e$ for each $t$, a contradiction. Hence, $C(T)$ is trivial.
Similarly, we can show that $C(T^{-1})$ is trivial. Thus $T$ acts distally on $G$. ∎
In view of Theorem 1.1 of [2], the following corollary is an easy consequence of Theorem 1.2. We will omit the proof.
@@
Let $G$ be a connected Lie group, $T\in{\rm Aut}(G)$ and let $L$ be a closed connected nilpotent normal $T$-invariant subgroup. Suppose $T$ acts trivially on both $L$ and $G/L$. If $T$ acts distally on ${\rm Sub}^{p}_{G}$, then $T={\rm Id}$.
@@
Step 1: If possible, suppose $T(x)\neq x$ for some $x\in U$. There exists a one-parameter subgroup $\{x_{t}\}$ in $G$ such that $x_{1}=x$. Let $H=\overline{\{x_{t}\}L}$. Then $H$ is connected and solvable, and $H/L$ is abelian. Let $\pi:G\to G/L$ be the natural projection. Then $\pi(H)=\overline{\pi(\{x_{t}\})}$. If $\pi(\{x_{t}\})$ is unbounded, it is closed and it is isomorphic to ${\mathbb{R}}$ and in this case $\{x_{t}\}$ is also unbounded and closed. Moreover $H=\{x_{t}\}L$.
Now suppose $\pi(\{x_{t}\})$ is relatively compact. As $H$ is solvable, there exists a maximal compact (connected, abelian) subgroup $K$ in $H$ such that $H=KL$. As $T(x)\neq x$ and $T|_{L}={\rm Id}$, we have that $T(h)\neq h$ for some $h\in K$, and hence $T(h)\neq h$ for some finite order element $h$ in $K$. There exists a closed (compact) one-parameter subgroup $\{h_{t}\}$ in $K$ such that $h_{1}=h$. We may replace $\{x_{t}\}$ by $\{h_{t}\}$ and assume also in this case that $H=\{x_{t}\}L$.
@@
as $n_{k}\to\infty$. Therefore, $y_{t}\in C$ for each $t$ and $\overline{\{y_{t}\}}\subset C$ as $C$ is closed. Also, $T^{n_{k}}(x_{t})=x_{t}y_{t}^{n_{k}}\to x_{t}y$ along a subsequence of $\{n_{k}\}$ for some $y\in\overline{\{y_{t}\}}\subset C$. Hence, $x_{t}\in C$, $t\in{\mathbb{R}}$. Thus $\{x_{t}\}\times\overline{\{y_{t}\}}\subset C$. Now $T(C)=C$, and we get that $T^{n_{k}}(T(\{x_{t}\}))\to C$. As $T$ acts distally on ${\rm Sub}^{p}_{G}$, we have that $T(\{x_{t}\})=\{x_{t}\}$. Since $T$ is unipotent, we have that $T(x_{t})=x_{t}$ for each $t$. But $T(x_{1})=x_{1}y_{1}\neq x_{1}$. This leads to a contradiction. Therefore, $T(x)=x$ for each $x\in U$, and hence, for each $x\in G$. Thus $T={\rm Id}$. ∎
@@
Step 2: Suppose $G$ is not nilpotent. Let $N$ be the nilradical of $G$. Then $T|_{N}={\rm Id}$. Suppose $G$ is solvable. As noted in Remark 3.3, we have that the automorphism induced by $T$ on $G/N$ is a finite order automorphism, and it is trivial, since it is unipotent. Now by Proposition 4.2, $T={\rm Id}$.
@@
We first assume that $G$ belongs to class ${\mathcal{C}^{\prime}}$, i.e.  $G$ has no compact subgroup of positive dimension. (For this case, we will follow the proof of Theorem 4.1 in [42] to some extent and then use Step 1 instead of Proposition 4.2 of [42].) Now ${\rm Aut}(G)$ is almost algebraic and it has finitely many connected components (see [12]). Thus some power of $T$ belongs to a one-parameter subgroup of ${\rm Aut}(G)^{0}$. Since $T$ is unipotent, we get that $T$ itself belongs to a one-parameter subgroup $\{T_{t}\}$ consisting of unipotent elements with $T=T_{1}$. Also, since $\{T^{n}\}_{n\in{\mathbb{Z}}}$ is co-compact in $\{T_{t}\}$, it is easy to see that each $T_{t}$ also acts distally on ${\rm Sub}^{p}_{G}$.
Let $G^{\prime}:=G/R$ and let $T^{\prime}_{t}$ be the automorphism of $G^{\prime}$ induced by $T_{t}$, $t\in{\mathbb{R}}$. Here, $T^{\prime}_{t}$ is also unipotent. Note that $G^{\prime}$ is semisimple and ${\rm Aut}(G^{\prime})^{0}$ is the group of inner automorphisms of $G^{\prime}$. There exists a one-parameter subgroup $\{u_{t}\}$ consisting of Ad-unipotent elements in $G^{\prime}$, such that $T^{\prime}_{t}={\rm inn}(u_{t})$, the inner automorphisms by $u_{t}$, $t\in{\mathbb{R}}$. As shown in [42] we can choose a Iwasawa decomposition $G^{\prime}=KAU$ such that $\{u_{t}\}\subset U$, and $U$ (resp. $AU$) is a closed simply connected nilpotent (resp. solvable) group. Let $\pi:G\to G^{\prime}$ denote the natural projection. Then $B:=\pi^{-1}(AU)$ is a closed connected solvable group, $T_{t}(B)=B$ and $T_{t}$ acts distally on ${\rm Sub}^{p}_{B}$. This implies from Step 2, that $T_{t}|_{B}={\rm Id}$. In particular, ${\rm Inn}(u_{t})$ acts trivially on $AU$, i.e. $\{u_{t}\}$ is contained in the center of $AU$. Let $\mathcal{A}$ (resp. $\mathcal{U}$) be the Lie algebra of $A$ (resp. $U$). Then $u_{t}=\exp tX$, $t\in{\mathbb{R}}$, for some $X\in\mathcal{U}$, and hence $X$ centralises $\mathcal{A}$. Using the restricted root system and Lemma 6.50 of [29], it follows that the centraliser of $\mathcal{A}$ in $\mathcal{A}\,\mathcal{U}$ is contained in $\mathcal{A}$. Therefore, $X\in\mathcal{A}\cap\mathcal{U}=\{0\}$. This implies that for each $t$, $u_{t}=e$, and hence that $T^{\prime}_{t}={\rm Id}$.
Let $G=SR$ be a Levi decomposition, where $S$ is a Levi subgroup (maximal connected semisimple subgroup) of $G$. Let $\pi_{1}:G\to G/N$ be the natural projection. Then $\pi_{1}(S)$ and $\pi_{1}(R)$ centralise each other. Since $T=T_{1}$ and $T^{\prime}_{1}={\rm Id}$ on $G/R$, we have that $T(x)\in xR$, $x\in G$. Let $x,y\in S$. Then $\pi_{1}(T(xyx^{-1}y^{-1}))=\pi_{1}(xyx^{-1}y^{-1})$. Since $[S,S]=S$, we get that $T(x)\in xN$ for all $x\in S$. As $N$ is closed, we get that $T(x)\in xN$ for all $x\in\overline{SN}$. As $T$ acts distally on ${\rm Sub}^{p}_{\overline{SN}}$ and $T|_{N}={\rm Id}$, by Proposition 4.2, we get that $T(x)=x$ for all $x\in S$. Since $T|_{R}={\rm Id}$. We get that $T={\rm Id}$.
We have shown that under the conditions stated in the hypothesis, $T={\rm Id}$ if $G$ is either solvable or if $G$ belongs to class ${\mathcal{C}^{\prime}}$.
Step 4: Let $M$ be the largest compact connected central subgroup of $G$. Suppose $M$ is non-trivial. Then $G/M$ belongs to class ${\mathcal{C}^{\prime}}$, $M$ is $T$-invariant and by Proposition 3.5, $T$ acts distally on ${\rm Sub}^{p}_{G/M}$. Since the automorphism of $G/M$ induced by $T$ is also unipotent, we get from Step 3 that $T$ acts trivially on $G/M$. That is, $T(x)\in xM$ for all $x\in G$. As $M$ is abelian, by Step 1, $T$ acts trivially on $M$. Now by Proposition 4.2, $T={\rm Id}$.
@@
Now we want to prove Theorem 1.4. We need an elementary result about certain automorphisms keeping a maximal compact connected abelian subgroup (maximal tori) invariant. In a connected Lie group $G$, all maximal tori are conjugate to each other. Any compact abelian subgroup of $G$ is contained in a maximal torus, as any maximal compact abelian subgroup of $G$ is connected. If $A$ is a maximal torus in $G$, then so is $T(A)$, for any $T\in{\rm Aut}(G)$, and $T(A)$ is conjugate to $A$. Any maximal compact subgroup of a connected solvable Lie group is abelian (cf. [25]).
@@
Let ${\mathcal{K}}$ be a compact connected abelian subgroup of ${\rm Aut}(G)$ and let $H={\mathcal{K}}\ltimes G$; where the action of ${\mathcal{K}}$ on $G$ is by automorphisms. Then $H$ is a connected Lie group and $G$ is normal in $H$. Let $R$ be the radical of $G$. Then $R$ is characteristic in $G$ and hence normal in $H$. Let $A$ be a maximal torus in $H$ containing ${\mathcal{K}}$. Arguing as above, we get that $A\cap G$ (resp.  $A\cap R$) is a maximal torus in $G$ (resp. $R$). As ${\mathcal{K}}\subset A$ and $A$ s abelian, we get that every automorphism in ${\mathcal{K}}$ acts trivially on $A\cap G$ (resp. $A\cap R$). ∎
@@
Let $G$ be a connected Lie group and let $T\in{\rm Aut}(G)$. The statements $(5)\implies(4)\implies(3)\implies(2)\implies(1)$ are obvious. Now it is enough to show that $(1)\implies(5)$ when $G$ is not a vector group and $T$ acts distally on the largest compact connected central subgroup (say) $M$ of $G$. Suppose (1) holds, i.e. $T$ acts distally on ${\rm Sub}^{p}_{G}$. Then we want to show that $T$ is contained in a compact subgroup of ${\rm Aut}(G)$. If $M$ is trivial, then $G$ belongs to the class ${\mathcal{C}^{\prime}}$ and the assertion follows from Theorem 1.1 (2). Now suppose $M$ is non-trivial. By Theorem 1.2, $T$ acts distally on $G$. Now $\overline{T}$ also acts distally on $G/M$ (cf. [34], Theorem 3.1).
@@
Now suppose $M$ is non-trivial and $G\neq M$. We show that there exists a maximal torus of $G$ which is $T^{k}$-invariant for some $k\in{\mathbb{N}}$. Let $\pi:G\to G/M$ be the natural projection. By Proposition 3.5, $\overline{T}$ acts distally on ${\rm Sub}^{p}_{G/M}$, where $\overline{T}$ is the automorphism of $G/M$ induced by $T$.
@@
Let ${\mathcal{K}}$ be the connected component of the identity in the closed subgroup generated by $\overline{T}$ in ${\rm Aut}(G/M)$, then $(\overline{T})^{k}\in{\mathcal{K}}$ for some $k\in{\mathbb{N}}$. Now by Lemma 4.3, $(\overline{T})^{k}$ keeps a maximal torus (say) $A$ of $G/M$ invariant and it acts trivially on $A$. Let $K=\pi^{-1}(A)$. Then $K$ is a maximal torus of $G$ and $T^{k}(K)=K$.
Note that $K$ is compact, connected and abelian. Moreover, as $T$ acts distally on $G$, so does $T^{k}$ and the restriction of $T^{k}$ to $K$ is also distal. Arguing as above for $K$ instead of $M$, we get that for some $m\in{\mathbb{N}}$, $T^{m}|_{K}$ is unipotent and, as it acts distally on ${\rm Sub}^{p}_{K}$, it acts trivially on $K$ (by Theorem 1.3). Without loss of any generality, we may replace $T$ by $T^{m}$ and assume that $T$ fixes every element of $K$, i.e. $T\in F_{K}(G):=\{\varrho\in{\rm Aut}(G)\mid\varrho(x)=x\mbox{ for all }x\in K\}$. We know that $F_{K}(G)$ is almost algebraic (cf. [15]). As $T$ acts distally on $G$ and $T\in F_{K}(G)$, for some $n\in{\mathbb{N}}$, $T^{n}$ is contained in a connected almost algebraic subgroup ×${\mathcal{B}}\times{\mathcal{U}}$, where ${\mathcal{B}}$ is a compact group and ${\mathcal{U}}$ consists of unipotent elements (cf. [1], Corollaries 2.3 and 2.5). Now we have that $T=\phi\circ S=S\circ\phi$ such that $\phi\in{\mathcal{B}}$ and $S\in{\mathcal{U}}$. As ${\mathcal{B}}$ is compact, we have that $S$ acts distally on ${\rm Sub}^{p}_{G}$. By Theorem 1.3, $S={\rm Id}$. Thus $T=\phi\in{\mathcal{B}}$ and (5) holds. ∎
@@
Let $G$ be a connected Lie group and let ${\mathcal{H}}$ be a subgroup of ${\rm Aut}(G)$. Suppose ${\mathcal{H}}$ acts distally on the largest compact connected central subgroup of $G$. Then so does $\overline{{\mathcal{H}}}$. Note that ${\mathcal{H}}$ acts distally on a compact ${\mathcal{H}}$-invariant subspace of ${\rm Sub}_{G}$ if and only if so does $\overline{{\mathcal{H}}}$ (both the above statements follow from Theorem 1 in [16]). Also (7) holds for ${\mathcal{H}}$ if and only if it holds when ${\mathcal{H}}$ is replaced by $\overline{{\mathcal{H}}}$, as ${\mathcal{K}}{\mathcal{D}}$ is closed in ${\rm GL}(d,{\mathbb{R}})$ for ${\mathcal{K}},{\mathcal{D}}$ as in (7). Therefore to prove the assertions, we may assume that ${\mathcal{H}}$ is closed. Note that every compact subgroup of ${\rm Aut}(G)$ acts distally on ${\rm Sub}_{G}$ (see e.g. [41], Lemma 2.2). Now $(6)\implies(5)\implies(4)\implies(3)\implies(2)\implies(1)$ are obvious.
First suppose that $G$ is not a vector group. We show that $(1)\implies(6)$. Suppose (1) holds, i.e. every $T\in{\mathcal{H}}$ acts distally on ${\rm Sub}^{p}_{G}$. By Theorem 1.4, every $T\in{\mathcal{H}}$ is contained in a compact subgroup of ${\rm Aut}(G)$. Recall that ${\rm Aut}(G)$ is identified with a closed subgroup of ${\rm GL}({\mathfrak{g}})$ via the map $T\to{\rm d\,}T$, where ${\mathfrak{g}}$ is the Lie algebra of $G$ and it is a $d$-dimensional vector space for some $d\in{\mathbb{N}}$ (cf. [9], see also [22]). Therefore, ${\mathcal{H}}$ is also a closed subgroup of ${\rm GL}({\mathfrak{g}})$ and the latter is isomorphic to ${\rm GL}(d,{\mathbb{R}})$, and we get by Theorem 1.1 of [17] (see also Proposition 2 of [39]) that ${\mathcal{H}}$ is compact. Thus (6) holds and $(1-6)$ are equivalent. Thus, if $G$ is not a vector group, the assertions in the theorem hold for this case.

### v3

@@ 4. Proofs of the main results
Let $\psi,S:H\to H$ be such that $S|_{K}={\rm Id}$, $S|_{C(T)}=\alpha\,{\rm Id}$, $\psi|_{K}={\rm Id}$ and $\psi|_{C(T)}=\phi$, where $\alpha$ and $\phi$ are as in Step 2. Now $T|_{H}=S\psi=\psi S$, and $\psi$ is contained in a compact subgroup of ${\rm Aut}(H)$. Therefore, $S$ acts distally on ${\rm Sub}^{p}_{H}$ (cf. [41], Lemma 2.2), and $C(S)=C(T)$.
@@
Let $\{z_{t}=x_{t}c_{t}\}$ be a closed one-parameter subgroup, where $\{x_{t}\}\subset K$ is closed and isomorphic to ${\mathbb{S}}^{1}$ and $\{c_{t}\}\subset C(S)=C(T)$ is a non-trivial (closed) one-parameter subgroup such that $S(c_{t})=c_{\alpha t}$, $t\in{\mathbb{R}}$ (as addition is the group operation on $C(T)={\mathbb{R}}^{d}$, we choose $c_{t}=tc$, $t\in{\mathbb{R}}$, for some nonzero $c\in{\mathbb{R}}^{d}$). Let $\{n_{k}\}$ be an unbounded sequence in ${\mathbb{N}}$ such that $S^{n_{k}}(\{z_{t}\})\to C$ for some $C$ in ${\rm Sub}^{p}_{H}$. Note that $\{z_{t}\}\subset\{x_{t}\}\times\{c_{t}\}$ and the latter is closed and $S$-invariant. So $C\subset\{x_{t}\}\times\{c_{t}\}$. For a fixed $t\in{\mathbb{R}}$, as $c_{t}\in C(S)=C(T)$, we get that
@@
As $\{x_{t}\}\times\{c_{t}\}$ is a direct product and $\{c_{t}\}$ is isomorphic to ${\mathbb{R}}$, it follows that $\beta=\alpha$ and $x_{t}=x_{\alpha t}$, i.e. $x_{(1-\alpha)t}=e$ for all $t\in{\mathbb{R}}$. As $\alpha\neq 1$, $x_{t}=e$ for each $t$, a contradiction. Therefore, $C(T)$ is trivial.
Similarly, we can show that $C(T^{-1})$ is trivial, and hence $T$ acts distally on $G$. ∎
In view of Theorem 1.1 of [2], the following corollary is an easy consequence of Theorem 1.2. We omit the proof.
@@
Let $G$ be a connected Lie group, $T\in{\rm Aut}(G)$ and let $L$ be a closed connected nilpotent normal $T$-invariant subgroup of $G$. Suppose $T$ acts trivially on both $L$ and $G/L$. If $T$ acts distally on ${\rm Sub}^{p}_{G}$, then $T={\rm Id}$.
@@
Step 1: If possible, suppose $T(x)\neq x$ for some $x\in U$. There exists a one-parameter subgroup $\{x_{t}\}$ in $G$ such that $x_{1}=x$. Let $H=\overline{\{x_{t}\}L}$. Then $H$ is a closed connected solvable $T$-invariant subgroup of $G$ and $H/L$ is abelian. Let $\pi:G\to G/L$ be the natural projection. Then $\pi(H)=\overline{\pi(\{x_{t}\})}$. If $\pi(\{x_{t}\})$ is unbounded, it is closed and it is isomorphic to ${\mathbb{R}}$ and in this case $\{x_{t}\}$ is also unbounded and closed. Moreover $H=\{x_{t}\}L$.
Now suppose $\pi(\{x_{t}\})$ is relatively compact. As $H$ is solvable, there exists a maximal compact (connected, abelian) subgroup $K$ in $H$ such that $H=KL$. As $T(x)\neq x$ and $T|_{L}={\rm Id}$, we have that $T(h)\neq h$ for some $h\in K$, and hence $T(h)\neq h$ for some finite order element $h$ in $K\cap U$. There exists a closed (compact) one-parameter subgroup $\{h_{t}\}$ in $K$ such that $h_{1}=h$. Replacing $\{x_{t}\}$ by $\{h_{t}\}$, we have in this case too that $H$ is closed and $H=\{x_{t}\}L$.
@@
as $n_{k}\to\infty$. Therefore, $y_{t}\in C$ for each $t$, and $\overline{\{y_{t}\}}\subset C$ as $C$ is closed. Also, $T^{n_{k}}(x_{t})=x_{t}y_{t}^{n_{k}}\to x_{t}y$ along a subsequence of $\{n_{k}\}$ for some $y\in\overline{\{y_{t}\}}\subset C$. Hence, $x_{t}\in C$, $t\in{\mathbb{R}}$. Thus $\{x_{t}\}\times\overline{\{y_{t}\}}\subset C$. Now $T(C)=C$, and we get that $T^{n_{k}}(T(\{x_{t}\}))\to C$. As $T$ acts distally on ${\rm Sub}^{p}_{G}$, we have that $T(\{x_{t}\})=\{x_{t}\}$, and since $T$ is unipotent, $T(x_{t})=x_{t}$ for each $t$. But $T(x_{1})=x_{1}y_{1}\neq x_{1}$. This leads to a contradiction. Therefore, $T(x)=x$ for each $x\in U$, and hence, for each $x\in G$. Thus $T={\rm Id}$. ∎
@@
Step 2: Suppose $G$ is not nilpotent. Let $N$ be the nilradical of $G$. As $T$ acts distally on ${\rm Sub}^{p}_{N}$, it follows from Step 1 that $T|_{N}={\rm Id}$. Suppose $G$ is solvable. As noted in Remark 3.3, we have that the automorphism induced by $T$ on $G/N$ is a finite order automorphism, and it is trivial since it is unipotent. Now by Proposition 4.2, $T={\rm Id}$.
@@
Now $G/N$ is reductive. Let $\pi:G\to G/N$ be the natural projection. Then $\pi(G)=SR^{\prime}$, an almost direct product of a semisimple Levi subgroup $S$ of $\pi(G)$ and the radical $R^{\prime}=\pi(R)$ of $\pi(G)$, where $R^{\prime}$ is central in $\pi(G)$. Let ${\overline{T}}$ be the automorphism of $\pi(G)$ induced by $T$. Then ${\overline{T}}$ acts trivially on $R^{\prime}$. Also ${\overline{T}}$ keeps $S$ invariant as it is the commutator subgroup of $\pi(G)$. Now if ${\overline{T}}$ acts trivially on $\pi(G)=G/N$, then by Proposition 4.2, $T={\rm Id}$. If possible suppose ${\overline{T}}$ does not act trivially on $\pi(G)$. As ${\overline{T}}$ is unipotent, there exists a nontrivial unipotent element $u$ in $S$ such that ${\overline{T}}={\rm inn}(u)$. Considering the Iwasawa decomposition $S=KAU$, where $u\in U$, we have that $u$ is not in the center of $AU$ (see Lemma 6.50 of [29]). Now $AUR^{\prime}$ is connected, solvable and ${\overline{T}}$ invariant. Let $B$ be the closure of $\pi^{-1}(AUR^{\prime})$. It is a closed connected solvable $T$-invariant subgroup of $G$. As $T$ acts distally on ${\rm Sub}^{p}_{B}$, by Step 2, $T|_{B}={\rm Id}$. This implies that ${\overline{T}}$ acts trivially on $AU$ and that $u$ centralises $AU$, a contradiction. Thus ${\overline{T}}$ acts trivially on $\pi(G)=G/N$, and as noted above, by Proposition 4.2, we get that $T={\rm Id}$.
@@
Now we want to prove Theorem 1.4. We need an elementary result about certain automorphisms keeping a maximal compact connected abelian subgroup (maximal torus) invariant. In a connected Lie group $G$, all maximal tori are conjugate to each other. Any compact abelian subgroup of $G$ is contained in a maximal torus, as any maximal compact abelian subgroup of $G$ is connected. If $A$ is a maximal torus in $G$, then so is $T(A)$, for any $T\in{\rm Aut}(G)$, and $T(A)$ is conjugate to $A$. Any maximal compact subgroup of a connected solvable Lie group is abelian (cf. [25]).
@@
Let ${\mathcal{K}}$ be a compact connected abelian subgroup of ${\rm Aut}(G)$ and let $H={\mathcal{K}}\ltimes G$; where the action of ${\mathcal{K}}$ on $G$ is by automorphisms. Then $H$ is a connected Lie group and $G$ is normal in $H$. Let $R$ be the radical of $G$. Then $R$ is characteristic in $G$ and hence normal in $H$. Let $A$ be a maximal torus in $H$ containing ${\mathcal{K}}$. Arguing as above, we get that $A\cap G$ (resp.  $A\cap R$) is a maximal torus in $G$ (resp. $R$). As ${\mathcal{K}}\subset A$ and $A$ is abelian, we get that every automorphism in ${\mathcal{K}}$ acts trivially on $A\cap G$ (resp. $A\cap R$). ∎
@@
Let $G$ be a connected Lie group and let $T\in{\rm Aut}(G)$. The statements $(5)\implies(4)\implies(3)\implies(2)\implies(1)$ are obvious. Now it is enough to show that $(1)\implies(5)$ when $G$ is not a vector group and $T$ acts distally on the largest compact connected central subgroup (say) $M$ of $G$. Suppose (1) holds, i.e. $T$ acts distally on ${\rm Sub}^{p}_{G}$. Then we want to show that $T$ is contained in a compact subgroup of ${\rm Aut}(G)$. If $M$ is trivial, then $G$ belongs to the class ${\mathcal{C}^{\prime}}$ and the assertion follows from Theorem 1.1 (2). Now suppose $M$ is non-trivial. By Theorem 1.2, $T$ acts distally on $G$. Then by Theorem 3.1 of [34], $\overline{T}$ acts distally on $G/M$, where $\overline{T}$ is the automorphism of $G/M$ induced by $T$.
@@
Now suppose $M$ is non-trivial and $G\neq M$. We show that there exists a maximal torus of $G$ which is $T^{k}$-invariant for some $k\in{\mathbb{N}}$. Let $\pi:G\to G/M$ be the natural projection. By Proposition 3.5, $\overline{T}$ acts distally on ${\rm Sub}^{p}_{G/M}$.
@@
Let ${\mathcal{K}}$ be the connected component of the identity in the closed (compact) subgroup generated by $\overline{T}$ in ${\rm Aut}(G/M)$. Then $(\overline{T})^{k}\in{\mathcal{K}}$ for some $k\in{\mathbb{N}}$. Now by Lemma 4.3, $(\overline{T})^{k}$ keeps a maximal torus (say) $A$ of $G/M$ invariant and it acts trivially on $A$. Let $K=\pi^{-1}(A)$. Then $K$ is a maximal torus of $G$ and $T^{k}(K)=K$.
Note that $K$ is compact, connected and abelian. Moreover, as $T$ acts distally on $G$, so does $T^{k}$ and the restriction of $T^{k}$ to $K$ is also distal. Arguing as above for $K$ instead of $M$, we get that for some $m\in{\mathbb{N}}$, $T^{m}|_{K}$ is unipotent and, as it acts distally on ${\rm Sub}^{p}_{K}$, it acts trivially on $K$ (by Theorem 1.3). Without loss of any generality, we may replace $T$ by $T^{m}$ and assume that $T$ fixes every element of $K$, i.e. $T\in F_{K}(G):=\{\varrho\in{\rm Aut}(G)\mid\varrho(x)=x\mbox{ for all }x\in K\}$. We know that $F_{K}(G)$ is almost algebraic (cf. [15]). As $T$ acts distally on $G$ and $T\in F_{K}(G)$, for some $n\in{\mathbb{N}}$, $T^{n}$ is contained in a connected almost algebraic subgroup ×${\mathcal{B}}\times{\mathcal{U}}$ of $F_{K}(G)$, where ${\mathcal{B}}$ is a compact group and ${\mathcal{U}}$ consists of unipotent elements (cf. [1], Corollaries 2.3 and 2.5). Now we have that $T=\phi\circ S=S\circ\phi$ such that $\phi\in{\mathcal{B}}$ and $S\in{\mathcal{U}}$. As ${\mathcal{B}}$ is compact, we have that $S$ acts distally on ${\rm Sub}^{p}_{G}$. By Theorem 1.3, $S={\rm Id}$. Then $T=\phi\in{\mathcal{B}}$ and (5) holds. ∎
@@
Let $G$ be a connected Lie group and let ${\mathcal{H}}$ be a subgroup of ${\rm Aut}(G)$. Suppose ${\mathcal{H}}$ acts distally on the largest compact connected central subgroup of $G$. Then so does $\overline{{\mathcal{H}}}$. Note that ${\mathcal{H}}$ acts distally on a compact ${\mathcal{H}}$-invariant subspace of ${\rm Sub}_{G}$ if and only if so does $\overline{{\mathcal{H}}}$ (both the above statements follow from Theorem 1 in [16]). Also (7) holds for ${\mathcal{H}}$ if and only if it holds when ${\mathcal{H}}$ is replaced by $\overline{{\mathcal{H}}}$, as ${\mathcal{K}}{\mathcal{D}}$ is closed in ${\rm GL}(d,{\mathbb{R}})$ for ${\mathcal{K}},{\mathcal{D}}$ as in (7). Therefore, to prove the assertions, we may assume that ${\mathcal{H}}$ is closed.
Note that every compact subgroup of ${\rm Aut}(G)$ acts distally on ${\rm Sub}_{G}$ (see e.g. [41], Lemma 2.2). Now $(6)\implies(5)\implies(4)\implies(3)\implies(2)\implies(1)$ are obvious.
First suppose that $G$ is not a vector group. We show that $(1)\implies(6)$. Suppose (1) holds, i.e. every $T\in{\mathcal{H}}$ acts distally on ${\rm Sub}^{p}_{G}$. By Theorem 1.4, every $T\in{\mathcal{H}}$ is contained in a compact subgroup of ${\rm Aut}(G)$. Recall that ${\rm Aut}(G)$ is identified with a closed subgroup of ${\rm GL}({\mathfrak{g}})$ via the map $T\mapsto{\rm d\,}T$, where ${\mathfrak{g}}$ is the Lie algebra of $G$ and it is a $d$-dimensional vector space for some $d\in{\mathbb{N}}$ (cf. [9], see also [22]). Therefore, ${\mathcal{H}}$ is also a closed subgroup of ${\rm GL}({\mathfrak{g}})$ and the latter is isomorphic to ${\rm GL}(d,{\mathbb{R}})$, and we get by Theorem 1.1 of [17] (see also Proposition 2 of [39]) that ${\mathcal{H}}$ is compact. Thus (6) holds and $(1-6)$ are equivalent. Thus, if $G$ is not a vector group, the assertions in the theorem hold for this case.